A survey was conducted to determine whether hours of sleep per night are independent of age. A sample of individuals was asked to indicate the number of hours of sleep per night with categorical options: fewer than 6 hours, 6 to 6.9 hours, 7 to 7.9 hours, and 8 hours or more. Later in the survey, the individuals were asked to indicate their age with categorical options: age 39 or younger and age 40 or older. sample data follow. Age Group Hours of Sleep 39 or younger 40 or older Fewer than 6 6 to 6.9 7 to 7.9 8 or more 38 62 75 65 36 59 73 92 (a) Conduct a test of independence to determine whether hours of sleep are independent of age. State the null and altemative hypotheses Ho: Hours of sleep per night is mutually exdusive from age. Ha: Hours of sleep per night is nat mutually exclusive from age. H: Hours of sleep per night is independent of age. Ha: Hours of sleep per night is not independent of age. H.: The proport on of people who get 8 or more hours of sleep per night is not equal across the two age groups. : The proportion of people who get 8 or more hours of sleep per night is equal across the two age groups. H: Hours of sleep per night is not independent of age. Ma: Hours of sleep per night is independent of age Find the value of the test statistic. (Round your answer to three decimal places.) What is the p-value? (Round your answer to four decimal places.) D-value Using a 0.05 level af significance, what is your condlusion? Reject H0. We conclude that hours of sleep per night and age are not independent. Do nat reject . We cennot conclude thet hours of sleep per night and age are not independent. O Do not reject H. We conclude thet hours of sleep per night and age are independent. o Reject Ho. We cannot conclude that hours of sleep per night and age are independent. (b) What is your estimate of the percentages of individuals who sleep fewer than 6 hours, to 6.9 hours, 7 to 7.9 hours, and 8 hours or more per night? Fewer than 6 6 to 6.9 7 to 7.9 8 or more
Answer #1

(a) right choice is second

H0:Hours of sleep per night is independent of age

Ha:Hours of sleep per night is not independent of age

test statistic=4.01

chi-square=sum((O-E)2/E)=4.01 with (r-1)(c-1)=(4-1)(2-1)=3 df

p-value=0.260

Donot reject H0: We conclude that hours of sleep per night and age are independent

(since the p-value=0.260 is more than level of significance =0.05)

(b) please look

%
<=6 14.8
6-6.9 24.2
7-7.9 29.6
>=8 31.4
total 100

following data is given

39 or younger 40 or older total %
<=6 38 36 74 14.8
6-6.9 62 59 121 24.2
7-7.9 75 73 148 29.6
>=8 65 92 157 31.4
total 240 260 500 100

here we use chi-square test and

chi-square=sum((O-E)2/E)=4.01 with (r-1)(c-1)=(4-1)(2-1)=3 df

Observed(O) Expected(E) E (O-E) (O-E)2/E
38 74*240/500 35.52 2.48 0.17
62 121*240/500 58.08 3.92 0.26
75 148*240/500 71.04 3.96 0.22
65 157*240/500 75.36 -10.36 1.42
36 74*260/500 38.48 -2.48 0.16
59 121*260/500 62.92 -3.92 0.24
73 148*260/500 76.96 -3.96 0.20
92 157*260/500 81.64 10.36 1.31
sum= 500 0 500 0 4.01
chi-square= 4.01 with 3 df
p-value= 0.2604